1.   Find the vector form of the equation of the plane that passes through the point P0 = (1, −2, 3) and has normal vector ~n =< 3, 1, −1 >.

2.   Find an equation for the plane that contains the line x = −1 + 3t, y = 5 + 3t, z = 2 + t and is parallel to the line of intersection of the planes x −2(y −1) + 3z = −1 and y = −2x −1 = 0.

3.   Find the point of intersection between the lines: < 3, −1, 2 > + t < 1, 1, −1 > and <−8, 2, 0 > + t < −3, 2, −7 >.

4.   Show that the lines x + 1 = 3t, y = 1, z + 5 = 2t for t ∈ R and x + 2 = s, y − 3 = −5s, z + 4 = −2s for t ∈ R intersect, and find the point of intersection.

5.   Find the point of intersection between the planes: −5x + y −2z = 3 and 2x −3y + 5z = −7.


1
Expert's answer
2021-07-13T13:16:55-0400

1. The standard form of the equation of the plane is


"3(x-1)+1(y-(-2))-1(z-3)=0"

"3x+y-z+2=0"

Parameters: "y=s, z=t"



"x=-\dfrac{2}{3}-\dfrac{1}{3}s+\dfrac{1}{3}t"

"y=0+1s+0t"

"z=0+0s+1t"

The vector form of the equation of the plane is


"\vec r=\langle-\dfrac{2}{3},0,0\rangle+s\langle-\dfrac{1}{3},1,0\rangle+t\langle\dfrac{1}{3},0,1\rangle"



2.


"x \u22122(y \u22121) + 3z = \u22121""y = \u22122x \u22121"




"y=-2x-1""x+4x+2+3z=-1"

"y=-2x-1""z=-\dfrac{5}{3}x-1""x\in \R"


The intersection between the planes is the line :


"\langle0,-1,-1\rangle+t\langle1, -2,-\dfrac{5}{3}\rangle, t\in \R"


The line


"\dfrac{x-0}{1}=\dfrac{y+1}{-2}=\dfrac{z+1}{-\dfrac{5}{3}}"


An equation for the plane parallel to the line of intersection of the planes


"ax+by+cz+d=0,"

where


"a-2b-\dfrac{5}{3}c=0"

The plane that contains the line "x = \u22121 + 3t, y = 5 + 3t, z = 2 + t"

"t=0: Point(-1, 5, 2)"

"t=-2: Point(-7, -1, 0)"

"-a+5b+2c+d=0""-7a-b+d=0""b=-7a+d""-a-35a+5d+2c+d=0""b=-7a+d""c=18a-3d"

Substitute


"a+14a-2d-30a+5d=0"

"d=5a"

If "a=1"

"d=5"

"b=-2"

"c=3"

The equation for the plane is


"x-2y+3z+5=0,"




3.


"3+s=-8-3t""-1+s=2+2t""2-s=0-7t""s=-3t-11""4=-5t-10""1=-5t+2"




"s=-3t-11""3=-12""1=-5t+2"

No solution.

There is no point of intersection.


4.


"3t-1=s-2""1=-5s+3""2t-5=-2s-4"




"s=3t+1""5s=2""2t=-2s+1"




"t=-\dfrac{1}{5}""s=\dfrac{2}{5}""t=\dfrac{1}{10}"


No solution.

There is no point of intersection.


5.


"-5x+y-2z=3""2x-3y+5z=-7"




"-13x-z=2""2x-3y+5z=-7"

"x=-\dfrac{2}{13}-\dfrac{1}{13}z"

"y=\dfrac{2}{3}x+\dfrac{5}{3}z+\dfrac{7}{3}"




"x=-\dfrac{2}{13}-\dfrac{1}{13}z"

"y=\dfrac{29}{13}+\dfrac{21}{13}z"

"z\in\R"

The intersection between the planes is the line :<3, - 1,2> +<1, 1, - 1>


"\langle-\dfrac{2}{13}, \dfrac{29}{13}, 0\rangle+t\langle-\dfrac{1}{13}, \dfrac{1}{13},1\rangle, t\in \R"